# Decision Synthesis in Monetary Policy

### v1

@@ Introduction
This is a model setting in which the decision variable is also modeled – it is an outcome variable. In the current quarter, $\mathbf{x}=(x_{1},\ldots,x_{k})^{\prime}$ is the $k-$vector of interest rate values over the next $k$ quarters, and $\mathbf{y}=(y_{1},\ldots,y_{k})^{\prime}$ the corresponding $k-$vector of inflation rates. Whatever other variables are in the VAR model ${\mathcal{M}}_{j},$ our interest focuses on the implied $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ required for BPDS eqn. (1). This conditional predictive is used in decision analysis with the same utility function for each model, namely $u_{j}(\mathbf{y},\mathbf{x})=U(\mathbf{y},\mathbf{x})$ given by
$U(\mathbf{y},\mathbf{x})=-\sum_{h=1{:}k}\{\rho^{k-h}(y_{h}-y^{*})^{2}+(x_{h}-x_{h-1})^{2}\}$ (5)
[…]
@@
The BPDS framework requires specification by the decision maker of a relevant class of baseline pdfs $p_{0}(\mathbf{y}|\mathbf{x}),$ the model-specific vector score functions $\mathbf{s}_{j}(\mathbf{y},\mathbf{x}),$ the initial BPDS model probabilities $\pi_{j}(\mathbf{x})$ as functions of candidate decisions $\mathbf{x},$ and the target expected scores $\mathbf{m}_{f}(\mathbf{x})$ at any $\mathbf{x}$. These are discussed in turn. In addition to customizing BPDS to the specific application, this section highlights a number of methodological developments relevant to other applications, particularly in: (a) the linkages of the $\pi_{j}(\mathbf{x})$ to $\mathbf{x}$ that are relevant more generally when $\mathbf{x}$ is both an outcome to be forecast as well as a putative decision variable (highlighted in Section 3.2.3 below); and (b) the relevance of dependence structure among the elements of the vector score under the initial distribution $p(\mathbf{y},{\mathcal{M}}_{j})$ (highlighted in Section 3.2.4 below).
@@
To complete the main BPDS pdf in eqn. (1) requires the baseline $p_{0}(\mathbf{y}|\mathbf{x}).$ This is taken as a multivariate T distribution with 10 degrees of freedom, using the location from the initial mixture $p(\mathbf{y}|\mathbf{x})$ ignoring the baseline (i.e., with $\pi_{0}(\mathbf{x})=0)$ and corresponding variance of that mixture inflated by 4. This defines a relevant, tractable ${\mathcal{M}}_{0}$ that can capture outcomes $\mathbf{y}$ that the two VAR models are not predicting well for any $\mathbf{x}$ under consideration, and signal that to the decision maker.
@@
The considerations of inflation targeting and interest rate smoothing reflected in the model-specific decision analysis in Section 3.1 are relevant to the choices of BPDS score functions. Our example takes $\mathbf{s}_{j}(\mathbf{y},\mathbf{x})=[s_{j1}(y_{1},x_{1}),\ldots,s_{jk}(y_{k},x_{k})]^{\prime}$ with elements
@@
where $y^{*}=2\%$ is the inflation target and $z_{y}$ and $z_{x}$ are score bandwidth parameters. This defines a class of bounded score functions, always relevant in decision analysis and here ensuring that the entropically tilted BPDS pdf of eqn. (3) is always integrable. The score bandwidths are set so that a certain deviation $d_{y}=(y-y^{*})^{2}$ has a score of ${\bm{\varepsilon}}$; given a choice of ${\bm{\varepsilon}}$ we set $z_{y}=d_{y}/\sqrt{-2\log({\bm{\varepsilon}})}$. Similar considerations apply to choosing $z_{x}$. Our analyses use ${\bm{\varepsilon}}=0.4$, $d_{y}=2$, and $d_{x}=1$ to ensure the score function is dispersed enough to accommodate modest changes in the Federal Funds rate while is more lenient in deviations from the inflation target. Obvious modifications could incorporate horizon $h-$specific inflation targets and differentially weight the two exponential terms, but this form suffices for our main goals in this paper. Note also that, if inflation deviations from target and interest rate changes are “small”, then $s_{jh}(y_{h},x_{h})$ is approximately quadratic in $|y_{h}-y^{*}|$ and $|x_{h}-x_{h-1}|$ for all $h,$ a perhaps more familiar utility form.
Initial Model Probabilities
For clarity in the presentation in this section, we now make explicit the dependency on time, so that the ingredients of the full BPDS predictive pdf in eqn. (1)– with the exponential form of the calibration function of eqn. (4)– are now indexed by current time $t$, i.e.,
@@
Bayesian model weighting based on historical predictive performance with respect to define forecast goals, as developed in (Lavine et al., 2021), provides the starting point for specification of the $\pi_{tj}(\mathbf{x}_{t}).$ The general form adopted is
@@
Traditional Bayesian analysis  (e.g. West and Harrison, 1997, chapter 12) defines the starting point. Here the time $t$ initial model probabilities are based on standard sequential Bayesian updating from those at $t-1$: That is, $\pi_{tj}\propto\pi_{t-1,j}p_{tj}(\mathbf{z}_{t-1,j}|{\mathcal{M}}_{j})$ where the “marginal model likelihood” term $p_{tj}(\mathbf{z}_{t-1,j}|{\mathcal{M}}_{j})$ is the value of the 1-step ahead predictive pdf under ${\mathcal{M}}_{j}$ at the observed values of the last period outcomes $\mathbf{z}_{t-1,j}$ under that model. In our applied setting, this $\mathbf{z}_{t-1,j}$ includes time $t-1$ outcome values of inflation $(y)$, interest rate $(x)$, and other economic indicators modelled and forecast in ${\mathcal{M}}_{j}$ in our setting. In general, these can differ across models, but in consideration for the initial weights we restrict to variables common across models.
Then, BPDS allows the decision maker freedom to make alternative choices of the $\pi_{tj},$ and the goal and decision focus recommends modification of the standard BMA choice. BMA, after all, only reweights models based on 1-step ahead predictive accuracy. Hence we adopt two modifications, based on recent literature consonant with the goal foci.
First, we use simple power discounting of historically accrued support across models, in which the time $t-1$ to time $t$ evolution is reflected in $\pi_{tj}\propto\pi_{t-1,j}^{\gamma}p_{tj}(\mathbf{z}_{t-1,j}|{\mathcal{M}}_{j})$ where $\gamma$ is a discount factor in $(0,1]$, closer to 1 for most applications. This acts to discount historically accrued support for model $j$ at a per-time unit discount rate $\gamma$ prior to updating by the time $t-1$ information. Going back at least to Smith (1979) and then, in a formal dynamic model uncertainty context, West and Harrison (1989, chapter 12, p.445), power-discounting has been shown to be of value in empirical studies in implicitly allowing for time-variation in the predictive relevance of different models (e.g. Raftery et al., 2010; Koop and Korobilis, 2013; Zhao et al., 2016). Our case study below uses $\gamma=0.95$.
Second, reflecting the foci on specific predictive and decision goals, initial model probabilities should also generally be modified based on the recent relative performance of models with respect to the defined goals. This is the premise underlying the specific variants of BPS in the setting of adaptive variable selection, or BPS-AVS, in  Lavine et al. (2021), and related developments in Loaiza-Maya et al. (2021), for example. This leads to the immediate BPDS extension of these prior approaches in which the above reasoning is extended to define
@@
Here the discounted Bayesian model probabilities are further updated with AVS-style weights based on the the realized BPDS calibration function based on relative model scores based on the actual decision outcomes at the last time period. As a result, models are initially and naturally reweighted based on both predictive and decision outcome performance at the last time period.
In our applied setting this means that, as a result, models achieving “good” recent trajectories of interest rate smoothness, as well as relatively accurate forecasting performance of realized inflation outcomes, will be rewarded with higher initial BPDS model probabilities in looking forward to the next time point. And we note that the specification here can cut back to define special cases including BPS-AVS (by setting ${\bm{\tau}}_{t-1}(\mathbf{x}_{t-1})=\mathbf{0})$, and within that to traditional BMA (by setting $\gamma=1$) for comparisons.
Finally, the inclusion in BPDS of the baseline model and its forecast densities leads to a modification of these initial model probabilities to provide a non-zero value $\pi_{t0}$ for the baseline. We choose a fixed probability– in our analysis $\pi_{t0}=0.1$ at each time point $t$– and simply renormalize the $\pi_{tj}$ above over $j=1{:}J$ accordingly.
@@
As noted earlier, in our setting the future values of decision variables are also considered outcomes predicted under the models. The models each forecast the future evolution of interest rates as part of the complex, dynamic macroeconomic system, whereas for decisions we must condition on $\mathbf{x}_{t}$. This is reflected in the conditional (on $\mathbf{x}_{t}$) distributions $p_{tj}(\mathbf{y}_{t}|\mathbf{x}_{t},{\mathcal{M}}_{j})$ in BPDS where $\mathbf{x}_{t}$ is treated as known. The theoretical implication for the BPDS model probabilities is the term $p_{tj}(\mathbf{x}_{t}|{\mathcal{M}}_{j})$ in eqn. (7)– this is the value of the current marginal predictive pdf of the vector $\mathbf{x}_{t}$ under ${\mathcal{M}}_{j}.$ Assuming the prior (to time $t$) probabilities $\pi_{tj}$ are specified, this form arises directly via Bayes’ theorem; the act of conditioning on $\mathbf{x}_{t}$ is informative, and the implied update is, simply by Bayes’ theorem, that in eqn. (7). Critically, this implies that candidate decision values that are not well-supported under the joint distribution of a model are down-weighted. Conversely, at any candidate decision vector $\mathbf{x}_{t},$ models that are more predictively supportive of the decision $\mathbf{x}_{t}$ will be relatively rewarded with higher values of resulting $\pi_{tj}(\mathbf{x}_{t}).$
In other applications of BPDS the decision variables may be exogenous, i.e., control variables that are to be chosen by the decision maker but that are not forecast jointly with $\mathbf{y}_{t}$ in the set of models. In such cases, it will be common to assume that the external choice of $\mathbf{x}_{t}$ is not informative, and then eqn. (7) results in decision-independent BPDS probabilities $\pi_{tj}(\mathbf{x}_{t})=\pi_{tj}$ based only on historical data and information.
@@
The BPDS target expected score $\mathbf{m}_{f}(\mathbf{x})=E_{f}[\mathbf{s}(\mathbf{y},\mathbf{x})]$ represents a desired improvement over the initial expected score $\mathbf{m}_{p}(\mathbf{x})=E_{p}[\mathbf{s}(\mathbf{y},\mathbf{x})]$. In the multi-objective case the resulting ${\bm{\tau}}(\mathbf{x})$ that defines $f(\mathbf{y}|\mathbf{x})$ to satisfy this target expectation is sensitive to both the relative scales and dependence of elements of $\mathbf{s}(\mathbf{y},\mathbf{x})$ under the initial mixture $\mathbf{y}\sim p(\mathbf{y}|\mathbf{x})$ at any candidate decision $\mathbf{x}.$ As functions of $\mathbf{y}$ the elements of the random score vector $\mathbf{s}(\mathbf{y},\mathbf{x})$ can be strongly correlated, leading to challenges in specifying relevant targets. This can also complicate the calculation of the implied BPDS tilting vector ${\bm{\tau}}(\mathbf{x})$– the vector that is needed to satisfy $\mathbf{m}_{f}(\mathbf{x})=E_{f}[\mathbf{s}(\mathbf{y},\mathbf{x})]$ under the BPDS density of eqn. (1). We address this by explicitly recognizing score dependencies and defining an approach that explicitly incorporates dependence.
@@
Some theoretical intuition is gained by considered cases of “small perturbations” in which $\mathbf{m}_{f}(\mathbf{x})-\mathbf{m}_{p}(\mathbf{x})$ has small elements. In this setting, entropic tilting theory in Tallman and West (2022) yields the second-order approximation ${\bm{\tau}}(\mathbf{x})\approx\mathbf{V}_{p}(\mathbf{x})^{-1}(\mathbf{m}_{f}(\mathbf{x})-\mathbf{m}_{p}(\mathbf{x}))$ where $\mathbf{V}_{p}(\mathbf{x})$ is the variance matrix of $\mathbf{s}(\mathbf{y},\mathbf{x})$ under the initial mixture $p(\mathbf{y}|\mathbf{x}).$ This shows that the implied tilting vector will be very sensitive to the initial score scales and dependencies as reflected in $\mathbf{V}_{p}(\mathbf{x}),$ and suggests a prime focus on a standardized score scale. That is, define $\mathbf{C}_{p}(\mathbf{x})$ as the scaled eigenvector matrix such that $\mathbf{V}_{p}(\mathbf{x})=\mathbf{C}_{p}(\mathbf{x})\mathbf{C}_{p}(\mathbf{x})^{\prime}$ and set the target score using $\mathbf{m}_{f}(\mathbf{x})=\mathbf{m}_{p}(\mathbf{x})+\mathbf{C}_{p}(\mathbf{x}){\bm{\epsilon}}(\mathbf{x})$ for a specified standardized expected score vector ${\bm{\epsilon}}(\mathbf{x})$. The usual convention is taken in which the eigenvector columns of $\mathbf{C}_{p}(\mathbf{x})$ are ordered according to decreasingly values of the corresponding eigenvalues, so that the first column is “dominant”, and so forth. This provides insights into how to practically define target scores related to the absolute standardized scale. As examples of the two extremes, taking ${\bm{\epsilon}}(\mathbf{x})=\epsilon(\mathbf{x})\mathbf{1}$ for some scalar $\epsilon(\mathbf{x})$ represents targets deviating from the initial expected score in equal amounts of $\epsilon(\mathbf{x})$ along each of the standardized eigendimensions. At the other extreme, and most relevant when there are strong score dependencies, taking ${\bm{\epsilon}}(\mathbf{x})=(\epsilon(\mathbf{x}),0,\ldots,0)^{\prime}$ defines the resulting target $\mathbf{m}_{f}(\mathbf{x})$ based on the major, dominant eigen-dimension alone. The latter is a starting point in general and is taken to define our BPDS case study below. In that setting, we choose $\epsilon(\mathbf{x})$ such that $\min\{(\mathbf{m}_{f}(\mathbf{x})/\mathbf{m}_{p}(\mathbf{x}))\}=0.75$ to define the maximum expected improved score in any dimension. [^2] Additionally, due to the arbitrariness of signs of eigenvectors, we apply a ±$\pm 1$ multiplier to the first column of $\mathbf{C}_{p}(\mathbf{x})$ so that the sum of elements are positive, ensuring the target score improves upon $\mathbf{m}_{p}(\mathbf{x})$. It is obviously straightforward to extend this methodology to define target scores impacted by higher eigen-dimensions, though that is left for future applications.
@@
The final step in BPDS is to couple the decision maker’s utility function with the BPDS predictive equations (1,2) to define the optimal decisions from the model synthesis. The decision maker can adopt any utility function but an initial neutral analysis will be based on using the same form as usual in the model-specific decisions, the function $U(\mathbf{y},\mathbf{x})$ of eqn. (5). This is used in the example analysis to follow, with the aim of computing $\mathbf{x}$ to maximize the implied expected utility function $\bar{U}(\mathbf{x}).$ In the case study analysis we compare decisions recommended by BPDS to those from each of the models and to a traditional BMA-based analysis. On the latter, the BMA mixture uses model weights proportional to the marginal likelihoods of the data which is common to all of the models (which includes inflation, interest rate, and GDP) under each ${\mathcal{M}}_{j}$. The BMA mixture naturally involves only the $p_{j}(\mathbf{y}|\mathbf{x})$ with no BPS/BPDS outcome-dependence, no notion of a baseline model to address model-set incompleteness, and no regard for the decision-focused use of the models. The foundational BPDS framework theoretically allows for these critical considerations as fundamental to the broader subjective Bayesian decision-analytic and goal-focused approach. Then, technically, BMA arises as a special case of the BPDS analysis as earlier discussed throughout Section 3.
@@
The computation of BPDS involves two key components. First, is the overall optimization over $\mathbf{x}$ which explores potential BPDS decisions and finds the optimizing vector $\mathbf{x}$. This requires an “outer loop” numerical optimization to explore $\mathbf{x}$ space. In our study, this is performed using a trust region method, namely Powell’s Derivative Free Optimization Solvers (PDFO–Ragonneau and Zhang, 2023). Due to the possibility of multi-modality in $\bar{U}(\mathbf{x})$, the optimization is run repeatedly (in parallel) from multiple starting values. In some periods over time, we do find evidence of multi-modality, so repeat starts of the optimization routine are mandated. Second, within each evaluation of a potential BPDS decision, it is necessary to compute the tilting vector ${\bm{\tau}}(\mathbf{x})$ given the constraint $E_{f}[\mathbf{s}(\mathbf{y},\mathbf{x})]=\mathbf{m}_{f}(\mathbf{x})$ for a target expected score $\mathbf{m}_{f}(\mathbf{x}).$. The theoretical basis of this is an implicit equation that is solved via standard, generic numerical optimization methods. Relevant details follow Tallman and West (2023, section 4.4) and are summarized in our Appendix A.

### v2

@@ Introduction
To find optimal decisions requires choice of utility function. The form adopted here is as follows. In the current quarter, $\mathbf{x}=(x_{1},\ldots,x_{k})^{\prime}$ is the $k-$vector of interest rate values over the next $k$ quarters, and $\mathbf{y}=(y_{1},\ldots,y_{k})^{\prime}$ is the corresponding $k-$vector of inflation rates. Whatever other variables are in ${\mathcal{M}}_{j},$ interest focuses on the implied $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ required for BPDS eqn. (1). This conditional predictive is used in decision analysis with the same utility function for each model, namely $u_{j}(\mathbf{y},\mathbf{g},\mathbf{x})=U(\mathbf{y},\mathbf{g},\mathbf{x})$ given by
$U(\mathbf{y},\mathbf{g},\mathbf{x})=-\sum_{h=1{:}k}\{\theta(y_{h}-y^{*})^{2}+(1-\theta)(g_{h}-g^{*})^{2}+(x_{h}-x_{h-1})^{2}\}$ (5)
where $\mathbf{g}$ denotes GDP growth.
This is a conventional utility function that reflects the dual mandate of inflation rate targeting, moderate growth in real-activity, and interest rate smoothing over the next $k=8$ periods. The $\mathbf{y}$ terms relate an inflation targeting mandate of $y^{*}=2\%$ over the longer run. The $\mathbf{g}$ terms reflect GDP growth with a target of $g^{*}=2.5\%$, roughly the average growth rate of GDP from 1990 to 2024. The $\mathbf{x}$ terms encourage relatively constrained changes in quarter-to-quarter interest rates (a “don’t rock the boat” consideration, as large swings in interest rates can/will have otherwise undesirable effects on the macroeconomy). The model-specific optimal decision vector $\mathbf{x}_{j}$ then maximizes $E_{p_{j}}[U(\mathbf{y},\mathbf{g},\mathbf{x})|\mathbf{x}]$ over $\mathbf{x}.$ Again, this analysis is repeated each quarter over time, producing rolling updates of the “currently optimal” projections for interest rates over the coming eight quarters.
@@
BPDS requires specification of a relevant class of baseline pdfs $p_{0}(\mathbf{y}|\mathbf{x})$, the model-specific vector score functions $\mathbf{s}_{j}(\mathbf{y},\mathbf{g},\mathbf{x}),$ the initial BPDS model probabilities $\pi_{j}(\mathbf{x})$ as functions of candidate decisions $\mathbf{x},$ and the target expected scores $\mathbf{m}_{f}(\mathbf{x})$ at any $\mathbf{x}$. These are discussed in turn. Beyond customizing BPDS to the specific application, we highlight methodological developments relevant to other applications. This includes, in particular: (a) the linkages of the $\pi_{j}(\mathbf{x})$ to $\mathbf{x}$ that are relevant more generally when $\mathbf{x}$ is both an outcome to be forecast as well as a putative decision variable; see Section 3.2.3; (b) the relevance of dependence structure among the elements of the vector score under the initial distribution $p(\mathbf{y},{\mathcal{M}}_{j})$; see Section 3.2.4). Notation considers the target variable $\mathbf{y}$ as including both inflation and GDP unless specifically denoted (as in the score/utility functions).
@@
Completing the main BPDS p.d.f. in eqn. (1) requires the baseline $p_{0}(\mathbf{y}|\mathbf{x}).$ This is taken as a multivariate T distribution with 10 degrees of freedom, using the location from the initial mixture $p(\mathbf{y}|\mathbf{x})$, ignoring the baseline (i.e., with $\pi_{0}(\mathbf{x})=0)$ and the corresponding variance of that mixture inflated by 2. This defines a relevant, tractable ${\mathcal{M}}_{0}$ that can capture outcomes $\mathbf{y}$ that the set of VAR models are not predicting well for any $\mathbf{x}$ under consideration, and signal that to the decision-maker.
@@
The dual mandate and interest rate smoothing from the model-specific decision analysis in Section 3.1 are reflected in the BPDS score functions. We take
$s_{j}(\mathbf{y},\mathbf{g},\mathbf{x})=[s_{j1}(y_{1},x_{1}),s_{j1}(g_{1},x_{1})\ldots,s_{jk}(y_{k},x_{k})s_{jk}(g_{k},x_{k})]^{\prime}$ (6)
with elements
@@
$s_{jh}(g_{h},x_{h})=\exp{\{-(g_{h}-g^{*})^{2}/(2z_{g}^{2})}\}+\exp{\{-(x_{h}-x_{h-1})^{2}/(2z_{x}^{2})}\},\quad h=1{:}k,$ (8)
where $y^{*}=2\%$ is the inflation target, $g^{*}=2.5\%$ is the GDP growth target, and $z_{y}$, $z_{g}$ and $z_{x}$ are score bandwidth parameters. This defines a class of bounded score functions, always relevant in decision analysis and here ensuring that the entropically tilted BPDS p.d.f. of eqn. (3) is always integrable. The score bandwidths are set so that a certain deviation $d_{y}=(y-y^{*})^{2}$ has a score of ${\bm{\varepsilon}}$; given a choice of ${\bm{\varepsilon}}$, we set $z_{y}=d_{y}/\sqrt{-2\log({\bm{\varepsilon}})}$. Similar considerations apply to choosing $z_{x}$ and $z_{g}$. Our analyses use ${\bm{\varepsilon}}=0.4$, $d_{y}=2$, $d_{g}=2$, and $d_{x}=1$ to ensure the score function is dispersed enough to accommodate modest changes in the Federal Funds rate while being more lenient in deviations from the inflation target. Obvious modifications could incorporate horizon $h-$specific inflation targets and differentially weight the two exponential terms, but this form suffices for our main goals in this paper. Note also that, if inflation deviations from target and interest rate changes are “small,” then $s_{jh}(y_{h},x_{h})$ is approximately quadratic in $|y_{h}-y^{*}|$ and $|x_{h}-x_{h-1}|$ for all $h,$ perhaps a more familiar utility form.
Initial and Conditional Model Probabilities
For clarity in this section, we make explicit the dependency on time, so that the ingredients of the full BPDS predictive p.d.f. in eqn. (1)– with the exponential form of the calibration function of eqn. (4)– are now indexed by current time $t$; that is,
@@
Bayesian model weighting based on historical predictive performance with respect to defined forecast goals (Lavine et al., 2021) is the starting point for specification of the $\pi_{tj}(\mathbf{x}_{t}).$ The general form adopted is
@@
Traditional analysis  (e.g., West and Harrison, 1997, chapter 12) defines the time $t$ initial model probabilities as Bayesian updates from those at $t-1$; that is, $\pi_{tj}\propto\pi_{t-1,j}p_{tj}(\mathbf{z}_{t-1,j}|{\mathcal{M}}_{j})$ where the “model marginal likelihood” term $p_{tj}(\mathbf{z}_{t-1,j}|{\mathcal{M}}_{j})$ is the value of the one-step ahead predictive p.d.f. under ${\mathcal{M}}_{j}$ at the observed values of the last period outcomes $\mathbf{z}_{t-1,j}$ under that model. In our setting, this includes time $t-1$ outcomes of inflation $(y)$, interest rate $(x)$, and other economic indicators in ${\mathcal{M}}_{j}$. In general these can differ across models, but in consideration for the initial weights we restrict to variables common across models.
BPDS allows the decision-maker freedom to make alternative choices of the $\pi_{tj},$ and the goal and decision focus recommend modification of the standard BMA choice. BMA, after all, only reweights models based on one-step ahead predictive accuracy. Hence, we adopt two modifications based on recent literature consonant with the goal foci.
First, we use simple power discounting of historically accrued support across models, in which the time $t-1$ to time $t$ evolution is reflected in $\pi_{tj}\propto\pi_{t-1,j}^{\gamma}p_{tj}(\mathbf{z}_{t-1,j}|{\mathcal{M}}_{j})$, where $\gamma$ is a discount factor in $(0,1]$, closer to 1 for most applications. This acts to discount historically accrued support for model $j$ at a per-time unit discount rate $\gamma$ prior to updating by the time $t-1$ information. Going back at least to Smith (1979) and then, in a formal dynamic model uncertainty context, West and Harrison (1989, chapter 12, p. 445), power-discounting has been shown to be of value in empirical studies in implicitly allowing for time variation in the predictive relevance of different models (e.g., Raftery et al., 2010; Koop and Korobilis, 2013; Zhao et al., 2016). Our case study uses $\gamma=0.95$.
Second, initial model probabilities are modified based on the recent relative performance of models with respect to the defined goals. This is the premise underlying the specific variants of BPS in the setting of adaptive variable selection, or BPS-AVS, in  Lavine et al. (2021), and related developments in Loaiza-Maya et al. (2021), for example. This leads to the immediate BPDS extension of these prior approaches with
@@
Here the discounted past probabilities are updated with AVS-style weights using the realized BPDS calibration function with relative model scores based on the actual decision outcomes at the last time period. As a result, models are initially and naturally reweighted based on both predictive and decision outcome performance at the last time period.
The result at this stage is that models achieving “good” recent trajectories of interest rate smoothness, as well as relatively accurate forecasting performance of realized inflation outcomes, will be rewarded with higher initial BPDS model probabilities looking forward. Of course, the specification can cut back to define special cases including BPS-AVS (with ${\bm{\tau}}_{t-1}(\mathbf{x}_{t-1})=\mathbf{0})$, and within that to traditional BMA (with $\gamma=1$) for comparisons.
Finally, the inclusion in BPDS of the baseline model and its forecast densities leads to a modification of these initial model probabilities to provide a non-zero value $\pi_{t0}$ for the baseline. We choose a fixed probability– in our analysis $\pi_{t0}=0.1$ at each time point $t$– and simply renormalize the $\pi_{tj}$ over $j=1{:}J$ accordingly.
@@
As noted earlier, in our setting the future values of decision variables are also considered outcomes predicted under the models. The models each forecast the future evolution of interest rates as part of the complex, dynamic macroeconomic system, whereas for decisions we must condition on $\mathbf{x}_{t}$. This is reflected in the conditional (on $\mathbf{x}_{t}$) distributions $p_{tj}(\mathbf{y}_{t}|\mathbf{x}_{t},{\mathcal{M}}_{j})$ in BPDS where $\mathbf{x}_{t}$ is treated as known. The theoretical implication for the BPDS model probabilities is the term $p_{tj}(\mathbf{x}_{t}|{\mathcal{M}}_{j})$ in eqn. (9)– this is the value of the current marginal predictive p.d.f. of the vector $\mathbf{x}_{t}$ under ${\mathcal{M}}_{j}.$ Assuming the prior (to time $t$) probabilities $\pi_{tj}$ are specified, this form arises directly via Bayes’ theorem. The act of conditioning on $\mathbf{x}_{t}$ is informative, and the implied update is, simply by Bayes’ theorem, that in eqn. (9). Critically, this implies that candidate decision values that are not well-supported under the joint distribution of a model are down-weighted. Conversely, at any candidate decision vector $\mathbf{x}_{t},$ models that are more predictively supportive of the decision $\mathbf{x}_{t}$ will be relatively rewarded with higher values of resulting $\pi_{tj}(\mathbf{x}_{t})$.
This is keeping with Leeper and Zha (2003) who argue that reduced form models are appropriate to use for policy analysis if the policy does not deviate too much from past policy changes. If the decision is very unusual agents may think they are in a new policy regime and adjust their behavior rendering a reduced form model less useful (i.e. the Lucas Critique). The technique above addresses these concerns by explicitly increasing the weight on models where the policy decision represents only a “modest intervention’. Further, if the decision is unsupported by the model set then the baseline model, being dispersed with fat tails, will receive more weight. In this way BPDS up-weights models least susceptible to the Lucas Critique, and in periods when there is more uncertainty the weights increase on the baseline model.
In other applications of BPDS, the decision variables may be exogenous; that is, control variables that are to be chosen by the decision-maker but that are not forecast jointly with $\mathbf{y}_{t}$ in the set of models. In such cases, it will be common to assume that the external choice of $\mathbf{x}_{t}$ is not informative, and then eqn. (9) results in decision-independent BPDS probabilities $\pi_{tj}(\mathbf{x}_{t})=\pi_{tj}$ based only on historical data and information.
@@
The BPDS expected score $\mathbf{m}_{f}(\mathbf{x})=E_{f}[\mathbf{s}(\mathbf{y},\mathbf{x})]$ represents a target for improvement over the initial expected score $\mathbf{m}_{p}(\mathbf{x})=E_{p}[\mathbf{s}(\mathbf{y},\mathbf{x})]$. In the multi-objective case, the resulting ${\bm{\tau}}(\mathbf{x})$ that defines $f(\mathbf{y}|\mathbf{x})$ to satisfy this target expectation is sensitive to both the relative scales and dependence of elements of $\mathbf{s}(\mathbf{y},\mathbf{x})$ under the initial mixture $\mathbf{y}\sim p(\mathbf{y}|\mathbf{x})$ at any candidate decision $\mathbf{x}.$ As functions of $\mathbf{y}$, the elements of the random score vector $\mathbf{s}(\mathbf{y},\mathbf{x})$ can be strongly correlated, leading to challenges in specifying relevant targets. This can also complicate the calculation of the implied BPDS tilting vector ${\bm{\tau}}(\mathbf{x})$ (i.e., the vector that is needed to satisfy $\mathbf{m}_{f}(\mathbf{x})=E_{f}[\mathbf{s}(\mathbf{y},\mathbf{x})]$ under the BPDS density of eqn. (1)). We address this by explicitly recognizing score dependencies and defining an approach that explicitly incorporates dependence.
@@
Some theoretical intuition is gained by considered cases of “small perturbations” in which $\mathbf{m}_{f}(\mathbf{x})-\mathbf{m}_{p}(\mathbf{x})$ has small elements. In this setting, entropic tilting theory in Tallman and West (2022) yields the second-order approximation ${\bm{\tau}}(\mathbf{x})\approx\mathbf{V}_{p}(\mathbf{x})^{-1}(\mathbf{m}_{f}(\mathbf{x})-\mathbf{m}_{p}(\mathbf{x}))$ where $\mathbf{V}_{p}(\mathbf{x})$ is the variance matrix of $\mathbf{s}(\mathbf{y},\mathbf{x})$ under the initial mixture $p(\mathbf{y}|\mathbf{x}).$ This shows that the implied tilting vector will be very sensitive to the initial score scales and dependencies as reflected in $\mathbf{V}_{p}(\mathbf{x}),$ and suggests a prime focus on a standardized score scale; that is, define $\mathbf{C}_{p}(\mathbf{x})$ as the scaled eigenvector matrix such that $\mathbf{V}_{p}(\mathbf{x})=\mathbf{C}_{p}(\mathbf{x})\mathbf{C}_{p}(\mathbf{x})^{\prime}$ and set the target score using $\mathbf{m}_{f}(\mathbf{x})=\mathbf{m}_{p}(\mathbf{x})+\mathbf{C}_{p}(\mathbf{x}){\bm{\epsilon}}(\mathbf{x})$ for a specified standardized expected score vector ${\bm{\epsilon}}(\mathbf{x})$. The usual convention is taken in which the eigenvector columns of $\mathbf{C}_{p}(\mathbf{x})$ are ordered according to decreasingly values of the corresponding eigenvalues, so that the first column is “dominant,” and so forth. This provides insights into how to practically define target scores related to the absolute standardized scale. As examples of the two extremes, taking ${\bm{\epsilon}}(\mathbf{x})=\epsilon(\mathbf{x})\mathbf{1}$ for some scalar $\epsilon(\mathbf{x})$ represents targets deviating from the initial expected score in equal amounts of $\epsilon(\mathbf{x})$ along each of the standardized eigen dimensions. At the other extreme, and most relevant when there are strong score dependencies, taking ${\bm{\epsilon}}(\mathbf{x})=(\epsilon(\mathbf{x}),0,\ldots,0)^{\prime}$ defines the resulting target $\mathbf{m}_{f}(\mathbf{x})$ based on the major, dominant eigen dimension alone. The latter is a starting point in general and is taken to define our BPDS case study that follows. In that setting, we choose $\epsilon(\mathbf{x})$ such that $\min\{(\mathbf{m}_{f}(\mathbf{x})/\mathbf{m}_{p}(\mathbf{x}))\}=0.75$ to define the maximum expected improved score in any dimension. [^3] Additionally, due to the arbitrariness of the signs of eigenvectors, we apply a ±$\pm 1$ multiplier to the first column of $\mathbf{C}_{p}(\mathbf{x})$ so that the sum of the elements are positive, ensuring the target score improves upon $\mathbf{m}_{p}(\mathbf{x})$. It is obviously straightforward to extend this methodology to define target scores impacted by higher eigen dimensions, though that is left for future applications.
@@
The final step couples the decision-maker’s utility with the BPDS predictive eqns. (1,2) to define the optimal decisions from the model synthesis. The decision-maker can adopt any utility function, but an initial neutral analysis will be based on using the same form as usual in the model-specific decisions: the function $U(\mathbf{y},\mathbf{x})$ of eqn. (5). This is used here, computing $\mathbf{x}$ to maximize the implied expected utility function $\bar{U}(\mathbf{x}).$ We compare decisions recommended by BPDS to those from each of the models and to the traditional BMA-based analysis. BMA uses model weights proportional to the marginal likelihoods of the data that are common across models (including inflation, interest rate, and GDP). As earlier discussed in Section 3, BMA arises as a special case of the BPDS analysis.
@@
The computation of BPDS involves two key components. First, the overall optimization over $\mathbf{x}$ explores potential BPDS decisions and finds the optimizing vector $\mathbf{x}$. This requires an “outer loop” numerical optimization to explore $\mathbf{x}$ space. This is done using the MATLAB implementation of particle swarm optimization. We use this due to occasional multi-modality in $\bar{U}(\mathbf{x})$ and because the algorithm supports parallelization which significantly speeds up optimization. Model specific optimizations are done using a trust region method, namely, Powell’s Derivative Free Optimization Solvers (PDFO– Ragonneau and Zhang, 2023). Second, within each evaluation of a potential BPDS decision, it is necessary to compute the tilting vector ${\bm{\tau}}(\mathbf{x})$ given the constraint $E_{f}[\mathbf{s}(\mathbf{y},\mathbf{x})]=\mathbf{m}_{f}(\mathbf{x})$ for a target expected score $\mathbf{m}_{f}(\mathbf{x}).$ The theoretical basis of this is an implicit equation that is solved via standard, generic numerical optimization methods. Relevant details follow Tallman and West (2023, section 4.4) and are summarized in our Appendix A.